OpenStudy (anonymous):

The following function is one-to-one. Find its inverse. Find the domain and range of f and Please show all of your work. f(x) = (8+x)/(x)

OpenStudy (mathteacher1729):

How far do go into this problem before getting stuck?

OpenStudy (alfie):

\[y = \frac{8+x}{x}\] The inverse is to find everything in function of x.. so.. \[xy = 8+x \rightarrow xy - x = 8 \rightarrow x(y-1) = 8 \rightarrow x = \frac{8}{y-1}\]

OpenStudy (alfie):

Domain, denominator different from zero. So just.. \[x \neq 0\]

OpenStudy (anonymous):

alfie- how did you get this?

OpenStudy (alfie):

because your function has just one problem, the denominator. We can't divide by zero, so we have to assure that the denominator must be different from zero. Since you just have "x", you just say "x different from zero". About the inverse function, as I said, is just about calcs.

OpenStudy (anonymous):

ok, so, x \[\neq\] 0 is the answer?

OpenStudy (anonymous):

\[x \neq 0\]

OpenStudy (alfie):

\[f \ \ \ \ is \ \ \ \ defined \ \ \ \ \forall x \in \ R \neq 0\]

OpenStudy (alfie):

Or if you prefer... (-infinity,0) U (0; + infinity)

OpenStudy (anonymous):

ok- is that the complete answer to this equation then? Its asking for the inverse, domain and range.

OpenStudy (alfie):

inverse: \[f(y) = \frac{8}{y-1}\] Domain: x ∈ R , x ≠ 0. Range: R.

OpenStudy (anonymous):

Thank you Alfie! :)